# Efficiently Serving LLM Reasoning Programs with Certaindex

### v1

@@ Abstract
1 Introduction
Large language model (LLM) capabilities have advanced rapidly, evolving from basic language understanding to solving complex mathematical problems [5, 47, 38], generating and debugging code [20, 52, 53], and analyzing case law [25]. A key driver for this progress is the emergence of sophisticated inference-time algorithms that enhance LLM reasoning [48, 51, 16]. These “LLM reasoning” algorithms enable LLMs to reflect on their own outputs, explore different reasoning paths, and generate more robust and accurate answers to challenging questions. However, they also yield more output tokens, sometimes by orders of magnitude [38, 47, 9], driving up inference compute demands and response latencies.
Although reasoning algorithms are diverse, they share in common that they each admit a flexible trade-off between inference compute and output accuracy, known as inference-time scaling. For example, self-consistency (SC) [48] generates multiple solutions to a query and aggregates the final answer via majority voting. Higher accuracy might be achieved by generating more solutions, but at the cost of more GPU cycles. Similarly, Monte Carlo Tree Search (MCTS) [16, 12] expands nodes of a solution tree, each representing a possible solution step. Expanding more nodes into a deeper or wider tree increases the likelihood of finding a correct solution but consumes more GPU resources to decode additional tokens.
While reasoning algorithms and their inference-time scaling properties have been extensively studied in experimental settings, deploying them in real-world and multi-tenant environments raises additional challenges in resource allocation. First, different user queries may pose varying levels of difficulty to solve, resulting in different amounts of inference compute required. Second, assuming work-conservation, allocating more compute to one user query inevitably delays or reduces compute resources available for others. Third, the best end-to-end latency is achieved by allocating the least amount of time needed to reach a correct answer, and no more than that. All of these objectives call for adaptively allocating inference compute between different user queries to effectively balance accuracy and cost.
Figure 1: (a) Inference time scaling of self-consistency [48] on GSM8K [7]. As we uniformly increase the number of output tokens (x-axis) per question, the overall accuracy (y-axis) grows then plateaus (see §2.2). (b) Canonical resource allocation in existing systems vs. ideal allocation for four queries P1-P4 with vary difficulties. (c) Correlations between certaindex and the compute required to obtain a correct answer. Each point is a question. Statistically higher certaindex indicates lower compute needed.
Fig. 1(a) illustrates the characteristics of inference-time scaling on GSM8K [7], a popular math reasoning benchmark. While uniformly scaling inference compute (number of generated tokens) by 2.5×2.5$\times$2.5 × across all queries improves overall accuracy from 60% to 80%, several opportunities for better resource allocation remain. As shown in Fig. 1(b), harder queries (blue zone) such as P3 may not be allocated sufficient compute to reach the correct answer. Simpler queries (green zone) such as P1 could have completed faster with less compute. Other queries (e.g., P4) may remain unsolved regardless of compute, and could have been terminated earlier. An alternative allocation – e.g., shifting compute from P1 to P3 while decreasing compute for P4 – could have yielded both higher accuracy and lower total compute cost.
Although many LLM serving systems [23, 57, 1] exist, they all fall short in effectively scheduling and allocating inference compute for LLM reasoning programs. Their schedulers [13, 50] operate at the granularity of individual input/output sequences, while LLM reasoning algorithms may invoke multiple input/output sequences corresponding to, e.g., different paths in MCTS. This design yields control of inference-time scaling out of the scheduler entirely and into the hands of the user, who has only partial information about all active queries, or even the relative difficulty of their own query, which are crucial for effective compute allocation. Furthermore, existing systems are agnostic to which individual input/output sequences belong to a given parent LLM reasoning query, making it impossible for the scheduler to optimize for end-to-end latency SLOs of those parent queries. These shortcomings typically lead to naive or uniform inference compute scheduling that suffer from wasted compute, degraded accuracy, and missed latency targets.
To address these limitations, we designed Dynasor, a system that orchestrates inference-time compute for dynamically serving reasoning programs. Unlike prior inference serving systems, Dynasor is explicitly aware of, and can control, the schedule of input/output requests associated with each reasoning query. This enables Dynasor to measure the progress of different reasoning queries while they run, estimate their individual trade-offs between allocating more compute and progress towards a final answer. Dynasor adaptively allocates more compute for difficult queries, less compute for easy queries, and early-terminates out-of-reach queries. It also minimizes the end-to-end latency of reasoning queries by gang-scheduling their group of input/output requests to mitigate stragglers, and promotes fairness by moderating compute allocations across competing reasoning queries.
To measure a statistically meaningful notion of reasoning progress, Dynasor uses a novel proxy variable, which we refer to as certaindex, inspired by prior work on uncertainty quantification of LLMs [21, 22, 11, 15]. Intuitively, certaindex quantifies how certain the LLM is in approaching its final answer (which may be correct or incorrect) at reasoning. Fig. 1(c) plots the correlation between certaindex values and the amount of compute needed to yield correct answers. When certaindex is high, the LLM is certain about its current answer, suggesting that more compute would not change its answer and less compute can be allocated. When certaindex is low, the LLM is uncertain about its current answer, and allocating more compute to explore alternative solution paths may help it approach a more certain answer.
Crucially, while LLM reasoning algorithms are diverse, they typically all possess a notion of certaindex that can be efficiently estimated at inference-time (§3.1). For example, in self-consistency, certaindex can be computed as the entropy of answers generated across all solution trajectories. For algorithms that involve reward models, the outputs of the reward model itself can be used as a measure of certaindex. Dynasor uses certaindex as a key narrow interface between the system and its applications, allowing it to support a wide range of current and future reasoning algorithms.
We implement Dynasor as a scheduling layer compatible with existing serving engines [23, 24, 57]. Our evaluations on various datasets, LLMs, and reasoning algorithms show that in batch processing scenarios, Dynasor saves up to $50\%$ compute to reach the same overall accuracy; in online serving, it sustains up to $\times3$ more queries or 4.7×4.7$\times$4.7 × tighter latency SLOs at the same attainment rates.
Specifically, this paper makes the following contributions:
$\bullet$ We identify the trade-offs among inference-time compute, question difficulty, and task latency and accuracy in serving reasoning queries, and develop certaindex as an indicative signal for reasoning progress, and show it is simple to extract, effective, and general across diverse tasks and algorithms.
$\bullet$ We develop Dynasor, an adaptive inference-serving system that leverages certaindex as a key narrow interface between its scheduler and diverse reasoning applications. Dynasor adaptively allocates compute and schedules reasoning requests across queries, optimizing accuracy, cost, latency, and fairness.
$\bullet$ Comprehensive evaluations across various datasets and reasoning algorithms show that Dynasor significantly outperforms prior state-of-the-art systems, including SGLang [57] and Parrot [29], using up to 50% fewer tokens to reach similar overall accuracy in batch processing, and sustaining $\times3$ higher request arrival rates and 4.7×4.7$\times$4.7 × more stringent latency SLOs in online serving.
2 Background and Motivation
Traditional LLM services operate on a “plain token generation” paradigm: given an input prompt, the LLM generates output tokens sequentially until a termination condition is met, such as a stop token or maximum sequenece length, is met. More complex applications are implemented by composing workflows of individual input/output requests.
Today, LLMs are increasingly applied to solve harder tasks requiring logical reasoning, multi-step planning, and iterative workflows. This paradigm, known as LLM reasoning, often requires generating multiple output sequences to a user query, each representing a different reasoning trajectory, and responds with a final answer only after multiple reasoning trajectories have been generated and considered. This shift introduces a fundamental distinction between serving traditional LLM queries and reasoning programs.
Figure 2: Illustration of the workflow of different LLM reasoning algorithms discussed in §2.1
2.1 LLM Reasoning Algorithms
@@
Rebase. Rebase [51], as shown in Fig. 2(b), also begins by generating $n$ independent outputs $I_{1},I_{2},\dots,I_{n}$ from the input prompt $P$. Unlike SC, these outputs represent intermediate steps, which may or may not contain the final solution. They are ranked and assigned with scores $s_{1},s_{2},\dots,s_{n}$, typically by a learned reward model. Rebase selects nodes with higher normalized exponential scores $e^{s_{i}}/\sum_{j=1}^{n}e^{s_{j}}$ as parent nodes, from which the next $n$ reasoning steps are branched out. This process is repeated until $n$ candidate solutions are returned. The final result is aggregated by (weighted) majority voting across all $n$ candidate answers or selecting the top-scored answer (a.k.a. best-of-n). Rebase structures the reasoning process as a solution tree, where tree’s depth (i.e., solution steps to reach an answer) depends on LLM outputs and its width is exactly $n$, which controls its inference-time compute.
@@
Internalized Chain-of-thought (ICoT). Lastest LLMs, such as OpenAI o1 [38] and Qwen-QWQ [47], can internalize reasoning behaviors at training, without needing for explicit reasoning algorithms. These models generate extended chain-of-thoughts [49] sequences to derive candidate answers to reasoning queries. They iteratively refine these answers by reflecting on prior outputs and continuing sequential decoding until termination, yielding the final solution. The number of refinement rounds $n$, which corresponds to total decoded tokens, directly determines inference-time compute. We examine this approach further in §6.7.
2.2 Inference Time Scaling
In each algorithm above, a control parameter $n$, which we call knob, governs the inference-time compute and the number of output tokens generated. Regardless of their differences, they all share a flexible trade-off: increasing $n$ improves the likelihood of yielding the correct answer but requires more compute to generate additional tokens. Fig. 1 demonstrates this scaling behavior for SC on GSM8K: as the token budget per question increases (x-axis), overall accuracy rises (y-axis). Similar trends are observed for other algorithms and datasets (see §6) and are reported in recent LLM advancements, such as OpenAI o1 [38] and open models [47, 42].
The exact scaling curve depends on both dataset difficulty and model capabilities. Using Fig. 1 as an example, we identify three characteristics that influence resource allocation and scheduling in serving systems: (1) Easy Zone (green): Low token budgets suffice for simpler queries, and over-allocating tokens should be avoided to prevent unnecessary latency. (2) Scaling Zone (blue): This regime shows that increasing compute significantly boosts accuracy, highlighting the importance of efficient resource allocation to balance cost and accuracy. (3) Impossible Zone (gray): Additional computation provides no benefits on questions beyond the model’s capabilities. Early termination of queries in this zone can minimize unnecessary resource consumption.
2.3 Existing LLM Inference Systems
Prior works on serving systems have extensively optimized LLM inference at the system-level. Given independently arriving requests, these systems leverage request batching to maximize GPU utilization [55, 3, 18, 27], memory paging to mitigate memory fragmentation and enable higher concurrency [23], disaggregation of prompt and output token processing to prevent interference between requests in different phases [58, 39], and more. However, all of these works assume independent input/output requests, and are agnostic to higher-level applications, such as LLM reasoning programs, that may submit interdependent inference requests.
On the other hand, ParrotServe [29] and SGLang [57] co-design the frontend and backend of LLM inference servers, and are aware of multi-request applications. ParrotServe provides an abstraction for users to specify dependencies between requests, enabling more effective cross-request scheduling. SGLang introduces programming primitives for multi-request workflows, optimizing execution by reusing intermediate KV-cache memory across requests. Although both ParrotServe and SGLang address multi-request applications, they do not target LLM reasoning programs, which exhibit unique characteristics of inference-time scaling and statistical measures of progress, and demand a more adaptive method of scheduling offered by Dynasor.
Next, we develop the certaindex measure.
3 Certaindex Based Resource Allocation

### v2

# Efficiently Scaling LLM Reasoning with Certaindex

@@ 3.2 Effectiveness of Certaindex for Dynamic Token Budgeting in Reasoning Optimization
3.3 Dynasor: A Certaindex-Driven Adaptive Compute Scheduler
The previous section demonstrated how Certaindex effectively quantifies certainty during LLM reasoning, enabling dynamic token allocation and reducing waste. Building on this, we introduce Dynasor – a reasoning-aware, end-to-end serving system that leverages Certaindex for adaptive scheduling. In real-world deployments for LLM reasoning, a single user query often spawns multiple related sub-queries, which we refer to as “reasoning programs”. These programs consist of logically connected requests that share context and computation. Traditional schedulers treat these sub-queries independently, missing opportunities for optimization. Dynasor addresses this by introducing a lightweight, Certaindex-driven scheduler that optimally manages both individual requests and entire reasoning programs. Beyond simple early exits, Dynasor implements two other key mechanisms to accelerate inference:
Certaindex-Driven Adaptive Compute Allocation. Dynasor dynamically adjusts token budgets for individual requests based on their real-time Certaindex values. As validated in §3.2, a high Certaindex indicates that the answer has stabilized, allowing the scheduler to reduce the token budget and save resources. This fine-grained control minimizes waste on queries that have converged or are unproductive, enhancing overall efficiency.
Gang Scheduling. Reasoning programs usually generate multiple related requests in parallel or at differet stages. Gang scheduling batches or prioritizes these requests together to maximize shared computation (e.g., KV-cache utilization) or early-exit high-certaindex programs to free resources. This mechanism reduces latency and improves throughput by aligning resource allocation with real-time reasoning progress.
Implementation. Dynasor is implemented as a lightweight scheduling component in SGLang [35]. For each decoding step, Dynasor monitor certaindex to reallocate resources or terminate requests, and use gang scheduling to prioritize requests within a reasoning program. Dynasor only introduces $\sim$500 lines of code into the core system, and requires no change to the model weights, reasoning algorithms, or the core execution backend. It acts as a thin layer around the standard decoding loop, yielding substantial gains in latency, throughput, and overall token efficiency. We describe the detailed design and implementation of Dynasor in Appendix D.
3.4 Theoretical Basis for Early Exiting in CoT without Accuracy Loss
This section gives a sketch of the theoretical foundation for our early exiting strategy based on “Probe-In-The-Middle”. We show how it can terminate CoT reasoning without accuracy degradation, and we believe this theoretical grounding can be generalized to other reasoning algorithms. See Appendix E for a riguous proof.
Sketch of Proof. Our method assumes that the next-token distributions $P_{t}(Y_{t+1}|x,Y_{1..t})$ in a reasoning chain eventually converge to a stationary distribution $P_{*}(Y|x)$. Once $P_{t}=P_{*}$, further computational steps are redundant. The “Probe-In-The-Middle” technique empirically detects this convergence. If observed answers (and thus their empirical distributions $\hat{P}$) stabilize across several probing steps, it indicates the underlying true distributions $P_{t}$ (and their mixtures $\bar{P}$) have likely converged. To analyze this, we define mixture distributions:
Definition 1 (Mixture Distributions).
Consider $n$ samples, each corresponds to $P_{t},\dots,P_{t+n}$. Denote the mixture distribution of samples $i+1,\dots,i+k\in[1,\dots,n]$ to be $\bar{P}_{i}^{i+k}=\frac{1}{k}\sum_{j=1}^{k}P_{i+j}$.
Our formal justification shows that if our empirical stopping criterion is met, the empirical mixture distributions $\hat{P}$ are $\epsilon/3$-close (in Total Variation distance) to the true mixture distributions $\bar{P}$. This requires a sufficient number of probes $k$, as established by the following concentration bound:
Lemma 1.
If $k=\widetilde{\Omega}\left(\frac{M+\log(1/\delta)}{\epsilon^{2}}\right)$, then $\mathrm{TV}(\bar{P}_{l}^{l+t},\hat{P}_{l}^{l+t})\leq\epsilon/3$, for all $l=i+1,\dots,i+k$, and $t=k-1,k$, with $1-\delta$ probability, assuming we have $M$ disjoint groups of the outputs.
The empirical stopping criterion is triggered when the TV distance between consecutively estimated mixture distributions, $\hat{P}_{i}^{i+k}$ and $\hat{P}_{i+j}^{i+j+k}$ (and similarly for mixtures of size $k-1$), remains below a small threshold $\epsilon^{\prime}=\epsilon/3$ for all $1\leq j\leq k$ (or $11\leq j\leq k-$ 1 respectively). When this empirical stability is observed, and $k$ is sufficiently large as per Lemma 1 (ensuring $\mathrm{TV}(\bar{P},\hat{P})\leq\epsilon/3$), the triangle inequality implies that the true underlying mixture distributions $\bar{P}_{i}^{i+k}$ and $\bar{P}_{i+j}^{i+j+k}$ are also $\epsilon$-close to each other (i.e., $\mathrm{TV}(\bar{P}_{i}^{i+k},\bar{P}_{i+j}^{i+j+k})\leq\epsilon$).
The connection between the stability of these true mixture distributions and the stability of individual distributions $P_{t}$ (which ultimately implies convergence to $P_{*}$ given Assumption 1) is provided by:
Lemma 2.
If $\mathrm{TV}(\bar{P}_{i}^{i+k},\bar{P}_{i+j}^{i+j+k})=0$, for any $1\leq j\leq k$, and if $\mathrm{TV}(\bar{P}_{i}^{i+k-1},\bar{P}_{i+j}^{i+j+k-1})=0$, for any $11\leq j\leq k-$ 1, then Eq. (2) holds for any $t\geq i$, and $T<2k-1$.
Therefore, observing empirical stability with a sufficiently large $k$ indicates that the true individual distributions $P_{t}$ have stabilized and are $\epsilon$-close to the stationary distribution $P_{*}(Y|x)$. This justifies that early termination preserves accuracy up to a tolerance $\epsilon$. The effective $k$ must satisfy both the concentration requirement (Lemma 1) and be large enough relative to $k^{*}$ from Assumption 1.
4 Evaluation
To show the effectiveness of certaindex in real-world workloads, we evaluate certaindex and the end-to-end serving system Dynasor in CoT [16] and three other reasoning algorithms (SC [4], REBASE [13], MCTS [6; 5]) on diverse datasets [36; 37; 38; 29; 20; 39]. For brevity, we provide detailed experimental setup in Appendix F.
Our experiments try to answer two main questions: (1) Batch workload: How effectively does certaindex reduce token consumption while preserving accuracy in real-world batch inference workloads? (§4.1); (2) Online workload: To what extent does Dynasor improve end-to-end performance in online serving scenarios? (§4.2). In addition, we perform ablation studies in (§4.3) to compare certaindex against other signals and selection of certaindex thresholds at runtime. We also perform a few more Dynasor system ablation studies in Appendix G.
4.1 Batch Inference with Certaindex
To evaluate the effectiveness of certaindex in batch inference, we evaluate Dynasor on CoT with reasoning models, and compare the tokens-accuracy tradeoff curve for different setups. We also evaluate certaindex on other reasoning programs (SC, MCTS, REBASE) to see if the effect generalizes. See detailed experiment setup in Appendix F.1 for CoT, and Appendix F.2 for SC, MCTS and Rebase.
CoT. Figure 6 shows our approach achieve significant token savings across all configurations, reducing token usage by 11-29% while maintaining same accuracy as the baseline. For the 10% of problems where our method achieves the highest token reduction, we observe savings of 34% on AIME and 53% on MATH500. In particular, for the top 1% of problems, we achieve even more substantial reductions of 53% on AIME and 81% on MATH500. A similar result is shown in Deepseek-R1 (See Appendix G.1). The primary gains of the token savings comes from the effective probing mechanism to identify self-doubt in CoT reasoning, and using certaindex to early-stopping.
Figure 6: Comparing Dynasor Token-Accuracy Curve Across Deepseek Qwen-Distilled Model Scales (7B, 14B, 32B) and Datasets (AIME24, AMC23, Math500)
Figure 7: Token-to-accuracy metric on batch processing workloads. Mean performance and std (error bars) of 10 runs are reported. Baseline-even allocates resource uniformally across all reasoning program. Baseline-length uses Detect@knob (Table 1) as the program’s process signal.
SC, MCTS, Rebase. To show our method generalizes to other reasoning algorithms, we also compare Dynasor against a modified SGLang intra-program scheduler using the following policies (1) baseline-even, which allocates resources uniformly; and (2) baseline-length which uses Detect@knob, the cumulative tokens generated at a specific step (Table 3) as the proxy.
Figure 7 show certaindex consistently reduces token usage by 9–52% across workloads compared to both baselines, without loss in accuracy. Notably, we achieve over 47% savings on SC-GSM8K and over 50% on REBASE-Math, highlighting Dynasor’s efficiency across diverse benchmarks. These gains primarily come from the intra-program scheduling algorithm(§ D.2.1), which accurately identifies high-certaindex programs and terminates them early without accuracy loss. This early termination strategy significantly reduces resource consumption by eliminating unnecessary sampling compared with baseline-even. In contrast, baseline-length leads to accuracy degradation even with less aggressive compute pruning, highlighting the effectiveness of certaindex-based resource allocation.
4.2 Online Serving with Certaindex
We evaluate Dynasor against SGLang [35] and Parrot [40] with different inter-program schedulers. We use P90 deadline attainment as the SLO attainment (detailed setup in Appendix F.3). Each row in Figure 8 presents a key performance trade-offs in online serving: (a) Program arrival rate vs SLO attainment: what percentage of programs can meet the SLO as the arrival rate increases; (b) SLO scale vs SLO attainment: how tightly a service provider can configure the SLO while maintaining reliable attainment. (c) Tradeoff between Accuracy and SLO attainment.
Figure 8: Evaluation on 3 online workloads on Dynasor against baselines. Rows from top to bottom: (a) Program rate vs SLO attainment, (b) SLO scale vs SLO attainment, (c) Accuracy vs SLO Attainment.
(a) Rate vs. SLO attainment. Figure 8(a) shows that Dynasor achieves much higher sustainable request rates under P90 deadline attainment: 1.63.3×1.6-3.3$\times$1.6 - 3.3 × and 1.63.2×1.6-3.2$\times$1.6 - 3.2 × compared to SGLang and Parrot, resp. These gains stem from two key factors. First, our intra-program scheduler identifies and early-terminates programs with high confidence, freeing resources for pending requests. Second, our inter-program scheduler prioritizes requests from the same program, increasing turnover rate.
(b) SLO scale vs. SLO attainment. Figure 8(b) shows deadline attainment under a fixed request rate for different systems. At these rates, Dynasor allows tighter SLO scales: 1.34.7×1.3-4.7$\times$1.3 - 4.7 × tighter than SGLang, and 1.73.3×1.7-3.3$\times$1.7 - 3.3 × tighter than Parrot. The source of gain is similar to (a): intra-program scheduler early-terminates requests and increase turnover rate, both increases effective request rate.
(c) Accuracy vs SLO attainment. Figure 8 (c) (details in Appendix F.3) shows Dynasor achieving 0.7% - 2% higher accuracy than SGLang and Parrot across all three workloads at the same SLO attainment. This improvement results from Dynasor’s ability to redistribute compute between simple and hard queries, enabling it to solve more queries while maintaining SLOs that baselines could only match with significantly more compute.
Throughput. Our system shows equal average throughput (token per second) in all online settings compared to the baselines. The is because that under the given request rates, all workloads saturate the GPU memory and making the system memory bound. This also validates the introduced scheduler has no overhead.
Figure 10: Correlation between certainty measure and mean steps required on solvable problems.
4.3 Ablation Studies
Choosing Different Thresholds. We show that certaindex threshold selection is vital. Fig. 10 demonstrates the impact of different certaindex thresholds. For SC on GSM8k, we select an entropy threshold of 0.5, as higher thresholds led to overly aggressive termination and accuracy degradation. Similarly, for MCTS on ASDiv, we choose a score threshold of 0.4, as lower thresholds (<0.4) decreased accuracy while higher thresholds (>0.4) increased token consumption without proportional accuracy gains. These results highlight the importance of careful threshold selection in balancing compute efficiency and accuracy. In practice, Dynasor uses a profiler-guided approach to select these thresholds (see detailed discussion in Appendix D.2.2)
Choosing Different signals. One may ask if there exist other signals better than certaindex. We compare certaindex with two alternative metrics for estimating resource needs. (1) Reasoning path length: Longer reasoning paths (more tokens) often indicate harder problems, suggesting a potential correlation between path length and compute needs (e.g., more reasoning paths). (2) Mean normalized log probability: Established as a measure of LLM confidence [25], higher log probabilities may correlate fewer samples needed for correct answers in SC. Using the (SC, GSM8K, Llama3.1-8B-instruct) setup, we evaluated certaindex alongside these metrics. Four metrics were tested: certaindex’s entropy measure $\mathcal{H}$, mean output length [41], mean normalized log probability [25], and their linear combinations.
As shown in Fig. 10, $\mathcal{H}$ achieved the strongest correlation with ground truth compute requirements (Pearson Correlation of 0.68, Kendall’s Tau of 0.61), outperforming other metrics and matching complex combinations. These results confirm that certaindex is a simple yet effective proxy for estimating inference computational demands, offering robust performance across tasks and models. Our end-to-end token-to-accuracy evaluations in Appendix G compare different signals for resource allocation, confirming certaindex’s superior performance.
Other ablations. We also ablate the contribution to scheduling algorithm (§ G.2), fairness analysis (§ G.3), and comparing static thresholding vs fine-grained resource allocation (§ G.4).
5 Related Work
Prior works on LLM serving systems have advanced inference through system-level optimizations like batching [42; 43; 44; 45], memory paging [46], and PD disaggregation [47; 48], yet these typically treat requests as independent units. While more recent frameworks such as ParrotServe [40] and SGLang [35] address multi-request workflows by enabling dependency specification and KV-cache memory reuse, neither approach specifically targets LLM reasoning programs, which demand adaptive scheduling based on test-time scaling behaviors and statistical progress measures—capabilities uniquely provided by Dynasor. Concurrently, research into efficient LLM reasoning, particularly for long CoT processes suffering from excessive token generation, has explored techniques like model-merging [49], mid-generation self-evaluation [50], specialized finetuning [11; 51; 52; 12; 53; 54], and question difficulty estimation [55], with various reasoning workloads [56; 20; 39]. However, while these methods enhance the accuracy-computation trade-off, their reliance on model modifications or additional external components often complicates production deployment. Distinctly, our approach employs a lightweight proxy variable, enabling a non-invasive and efficient scheduling layer that seamlessly integrates with existing infrastructure.
6 Limitation and Broader Societal Impact Discussion
Limitation. Our study primarily focused on optimizing token allocation through certaindex, but did not explore its integration with advanced serving techniques like PD disaggregation or chunked prefill. The potential impact of these integrations on the latency-accuracy trade-off in real-world workloads remains an open area for future research.
Broader Impact. Improvements of Dynasor in LLM inference efficiency can further democratize access to powerful models. By reducing computational costs and energy consumption without model changes, our method enables more sustainable and scalable LLM serving. However, we acknowledge potential risks, including bias in early exits, security vulnerabilities through side-channel attack for certaindex in multi-tenant settings, and manipulation risks if malicious inputs exploit the mechanism.
7 Conclusion
In this paper, we show that LLM reasoning algorithms frequently exhibit answer stabilization, leading to unnecessary token generation. We developed certaindex, an algorithm-agnostic metric, to detect this stability and enable early exit from the reasoning process. When we implement Certaindex as a scheduler into Dynasor, our reasoning-aware serving system, this simple yet effective mechanism results in significant practical benefits, including up to 50% compute savings and 3.3x higher throughput in production systems, without accuracy drop. Certaindex thus represents a key advancement in making sophisticated LLM reasoning more efficient and scalable.
Appendix A Example of LLM Reasoning Algorithms
@@
Figure 11: Illustration of the workflow of different LLM reasoning algorithms discussed in Appendix A
@@
Rebase. Rebase [13], as shown in Fig. 11(b), also begins by generating $n$ independent outputs $I_{1},I_{2},\dots,I_{n}$ from the input prompt $P$. Unlike SC, these outputs represent intermediate steps, which may or may not contain the final solution. They are ranked and assigned with scores $s_{1},s_{2},\dots,s_{n}$, typically by a learned reward model. Rebase selects nodes with higher normalized exponential scores $e^{s_{i}}/\sum_{j=1}^{n}e^{s_{j}}$ as parent nodes, from which the next $n$ reasoning steps are branched out.
This process is repeated until $n$ candidate solutions are returned. The final result is aggregated by (weighted) majority voting across all $n$ candidate answers or selecting the top-scored answer (a.k.a. best-of-n). Rebase structures the reasoning process as a solution tree, where tree’s depth (i.e., solution steps to reach an answer) depends on LLM outputs and its width is exactly $n$, which controls its inference-time compute.
@@
Internalized Chain-of-thought (ICoT). Lastest LLMs, such as OpenAI o1 [15] and Qwen-QWQ [34], can internalize reasoning behaviors at training, without needing for explicit reasoning algorithms. These models generate extended chain-of-thoughts [16] sequences to derive candidate answers to reasoning queries. They iteratively refine these answers by reflecting on prior outputs and continuing sequential decoding until termination, yielding the final solution. The number of refinement rounds $n$, which corresponds to total decoded tokens, directly determines inference-time compute.
Appendix B Certaindex Based Resource Allocation